OpenStudy (anonymous):

If f(x) = −2x − 4 and g(x) = x + 4, find (f/g)(−8). −3 3 −1 1

hero (hero):

f(x) = -2x - 4 g(x) = x + 4 (f/g)(x) = f(x) ÷ g(x) = (-2x -4)/(x+4) (f/g)(-8) = f(-8) ÷ g(-8) = (-2(-8) -4)/(-8+4) = (16 - 4)/(-4) = (12/-4) = -3